# A Multi-Scale Cognitive Interaction Model of Instrument Operations at the Linac Coherent Light Source

### v1

@@ 3 Experiments and Measurements
An “experiment” here means the contiguous period—usually around 3-5 days—during which users are actively using an LCLS instrument to collect data. An experiment has many sub-steps. The sub-steps that we will be primarily concerned with are the contiguous periods of time, called “measurements”, during which the x-ray beam is striking a given target—or at least intended to be doing so—with the intention of collecting data about that target under given conditions. [^3] Because the focused x-ray beam typically modifies the sample almost immediately, a given measurement typically utilizes a series of putatively identical targets, for example on a tape or embedded in a stream of liquid. We do not need to distinguish these cases, and so just use the term “target” or “sample” for either of these cases, unless it is necessary to distinguish them.
@@
Importantly, most experiments utilize a variety of targets, including “dummy” targets used to align and calibrate the instrument, “control” targets that have known properties, and “critical” targets for which the experiment is designed to gather new data. Because the critical targets are often costly and of limited availability, the experiment must be planned to maximize the high quality data—that is, data with low signal-to-noise ratio (SNR)—collected on these critical targets. Therefore, although an enormous amount of careful planning goes into experiment design well before the experimental team arrives at SLAC, unplanned issues always arise, so the team engages in continuous, on-site, real-time decision making to maximize the collection of high quality data on the critical targets.
The primary decisions to be made are how long a given measurement should continue, and in what order measurements should be taken from different targets. Because beam time is limited, it seems sensible to start gathering data on critical targets as soon as possible. However, because unexpected issues often arise during the experiment’s actual beam time, it may be more sensible to use dummy or control targets until it is clear that the instrument is working well (i.e., producing low SNR data), especially if the availability of critical targets is limited. On the other hand, as the experiment nears the end of its allocated beam time, the team may wish to get as much data as possible on critical targets, even if the instrument is not producing the expected level of SNR, and especially if things did not go as planned — for example due to unexpected linac down time — one may want to get as much data as possible on critical targets in the remaining time, regardless of the quality of the data being non-optimal. This sort of trade-off is very important, and, as will be seen in more detail later, defines the top layer of our model.
@@ 6.1 The Central Decision: What to Measure When and for How Long
However, recall that there may be limited sample material, especially of the most important samples, and the problems may resolve, improving the SNR. These considerations vie in the opposite direction, suggesting that if things are not working well (i.e., for whatever reason, the team is seeing low SNR in secondary data analysis), but it is still early in the overall experiment period, it may be better to hold on to the precious critical sample and hope that the problems resolve. But how long do you wait? Half way? Two thirds? Three quarters? And if the data is marginal, and getting worse, should you continue a measurement to wring as much out of the current sample, or simply abort it and move on to more important samples, given an impending hard stop? And once the ongoing measurement is completed, which of the 8x3x15 minute measurements should be tried next? Perhaps even a previous less-the-satisfactory measurement should be repeated?
@@ 7 A Multi-Scale Cognitive Interaction Model of LCLS Instrument Operations
Our goal in the present work is to predict the potential impacts of UI and workflow changes in the most impactful aspects of LCLS operations. Because of the difficulty of carrying out direct user studies, or of employing a “flight simulator” approach, in the LCLS setting, we chose to create a computer simulation (hereafter, “the model”) of some of the most important human-in-the-loop aspects of experiment operations, and in particular of the phenomena relevant to the core decision making, as discussed above, of when and whether to continue or abort a measurement, and what to measure next. The goal is to use this model to predict the direction and rough shape of changes in experiment efficiency that might result from changes in instrument procedures and/or control interfaces.
In what follows we discuss the central concepts represented in the model. Much more detail is included in the three video supplements to this paper[vids]. The first video is a concept review, and provides more-or-less the same information as below. The second includes several demonstrations of the model running, and discusses the results depicted in figures 7 through 9. The third video is a deep dive into the model code, which is available on github.[githubmodel].
@@ 7.2 The Micro-Cognitive Scale: Peak Chasing, Functional Acuity, and Functional Operability
Real experimental decision making takes place in a very complex setting, where many extrinsic factors are involved, such as which samples are available, what the value is of those samples, the stability and performance of all aspects of the system, how far through the experiment they are, and how much time the remaining high-value samples are likely to take, as well as large-scale factors such as SLAC and LCLS shift-changes and the requirements for relevant members of the team to take breaks to eat and sleep. Our model, of course, represents a highly simplified version of these, but it does not attempt to take into account all aspects of these factors.
@@ 7.3 The Meso-Cognitive Scale: When to End a Measurement
7.4 The Macro-Cognitive Scale: Sample Choice Planning and Replanning
The dynamic environment of an LCLS experiment requires a blend of meticulous pre-planning and continuous real-time, opportunistic decision-making and re-planning. This is markedly different from static, puzzle-like problem-solving where one has the luxury of time and a limited set of variables. The actual measurement process is akin to balancing a pencil on its tip, and is made more complicated by the nature of the novel physics being studied at LCLS, and, of course, by a constant barrage of tiny problems, requiring real-time decision-making and replanning. While some adjustments, like shortening a measurement if one is obtaining high-quality data, are relatively straightforward, others, such as extending a measurement, or even bringing in additional materials from off-site if necessary, can be more challenging given the strictly limited time frame of LCLS experiments, and are potentially impossible. In our model, the only information used to make these decisions is the information available regarding how well the measurements have been going so far, how much beamtime remains, and the number and quality of samples remaining to test.
@@ 7.5 Measurement Error and Performance Quality of Samples
Central to all of the above (and, of course, in the real world LCLS setting) is the quality of the data being gathered. Data quality in settings such as LCLS is often measured in terms of signal-to-noise ratio (SNR), as variously mentioned above. SNR is the slope of a scatter plot of the data, divided by its standard error, and standard error is simply the standard deviation of the data divided by the square root of the number of data points. Therefore, in our model we simply use standard error as a stand-in for SNR, and the goal of a given measurement is to reach a particular target standard error, which we call “Target Error” (TE), and which is 0.001 by default, for its given sample.
@@ 8 Example Results
Figure 7 depicts two explorations with the model. The left graph in Figure 7 scans across three values of Functional Acuity (FA), and is made with Adjust-Error=False and Cutoff-time=False, thus allowing the user to remain as long as necessary to reach the default Target Error (TE) of 0.001. Recall that runs start at high Performance Quality (PQ) and go to progressively lower PQs — that is, from lower importance samples to higher, or right to left on these graphs. Unsurprisingly, it requires less data to reach the TE threshold (0.001) as PQ increases, and as FA improves (becomes smaller), so that the operator notices that the beam and stream are not aligned at smaller distances of misalignment. The right graph in Figure 7 holds FA to 0.1 (the best performing value from the left), and compares Noticing Delays (NDs). In this case, Adjust-Error=True and Cutoff-time=True, which is the most realistic case: The team is going to get kicked off the machine at a fixed time, but they are allowed to adjust the Error Target to try to get some data on all samples. At first glance, in this case, Noticing Delay makes no apparent difference; All the lines are more-or-less the same. However, as PQ increases, the amount of data gathered appears to be smaller at both high and low PQ, but higher at middle values of PQ. The reason that this happens is that the experiment manager observes that they are going to run out of time with the default TE, and adjusts the TE upward so that some data is obtained for even the last/most important/lowest PQ samples, although they do run out of time in the end anyway. Again, more flexible re-planning would improve the data quality for the most important (lowest PQ) samples, as will be explored in a separate paper.
@@ 9 Discussion
Modulating Functional Acuity (FA) and Noticing Delay (ND) highlights how improved FA (i.e., faster response to beam mis-alignment) leads to increased overall experimental efficiency with fewer samples needed for success. In contrast, improved ND produces less significant impacts on efficiency, except in unusual operating regimes of the system, for example when there is no beam time limit. We also investigated quality and duration trade-offs by adjusting whether the target SNR (Target Error) is dynamic or fixed, and by imposing cutoff times, or allowing runs to continue indefinitely. Unsurprisingly, fixed target error requirements produce higher fidelity given unlimited time. However, allowing the target error to adapt yields better time management, ensuring sufficient yields from high-performing samples even if lesser-quality measurements become truncated by virtue of limited beam time.
@@
Our model merely scratches the surface of the potential of model-based cognitive engineering efforts as applied to scientific user facilities such as LCLS. The present model requires further elaboration and validation with respect to detailed measurements of individual and team behavior. It should be extended to capture more inter- and intra-team communications. Perhaps most importantly, models of learning at all levels should be introduced in order to capture both short-term individual learning, for example the instrument scientists becoming familiar with each new experiment; social learning, e.g., the team learning to work together; and longer-term learning, such as staff training.
As science probes deeper into the complexities of nature, large-scale, beyond-the-state-of-the-art efforts like those at LCLS, and other scientific user facilities, are rapidly increasing in importance. These facilities are in constant flux at every scale. Efforts such as ours that attempt to model the system at multiple scales may be helpful in increasing the efficiency of the science done at these sites by helping to smooth and optimize the human-in-the-loop workflow that will undoubtedly be a significant component of these operations of scientific user facilities for the foreseeable future.
@@ 10 Acknowledgements
The research described here arises from a project begun in 2019 by Devangi Vivrekar and Paul Fuoss, with the assistance of Jeff Shrager. The research was picked up by Teddy Rendahl, and then by Wan-Lin Hu and Jonathan Segal. The numerous LCLS and SLAC staff and visiting scientists that permitted us to observe them working, and sat for interviews over those years, are too numerous to mention, but we are especially grateful to Andy Aquila, Mark Hunter, Meng Liang who welcomed us into their experiments, and to Dan Flath and especially Jana Thayer for valuable discussions and support. This project was supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences under Contract No. DE-AC02-76SF00515.

### v2

@@ 3 Experiments and Measurements
An “experiment” here means the contiguous period, usually around 3-5 days, during which users are actively using an LCLS instrument to collect data. An experiment has many sub-steps. The sub-steps that we will be primarily concerned with are the contiguous periods of time, called “measurements”, during which the x-ray beam is striking a given target—or at least intended to be doing so—with the intention of collecting data about that target under given conditions. [^3] Because the focused x-ray beam typically modifies the sample almost immediately, a given measurement typically utilizes a series of putatively identical targets, for example on a tape or embedded in a stream of liquid. We do not need to distinguish these cases, and so just use the term “target” or “sample” for either of these cases, unless it is necessary to distinguish them.
@@
Importantly, most experiments utilize a variety of targets, including “dummy” targets used to align and calibrate the instrument, “control” targets that have known properties, and “critical” targets for which the experiment is designed to gather new data. Because the critical targets are often costly and of limited availability, the experiment must be planned to maximize the high quality data — that is, data with high signal-to-noise ratio (SNR) — collected on these critical targets. Therefore, although an enormous amount of careful planning goes into experiment design well before the experimental team arrives at SLAC, unplanned issues always arise, so the team engages in continuous, on-site, real-time decision-making to maximize the collection of high quality data on the critical targets.
The primary decisions to be made are how long a given measurement should continue, and in what order measurements should be taken from different targets. Because beam time is limited, it seems sensible to start gathering data on critical targets as soon as possible. However, because unexpected issues often arise during the experiment’s actual beam time, it may be more sensible to use dummy or control targets until it is clear that the instrument is working well (i.e., producing high SNR data), especially if the availability of critical targets is limited. On the other hand, as the experiment nears the end of its allocated beam time, the team may wish to get as much data as possible on critical targets, even if the instrument is not producing the expected level of SNR, and especially if things did not go as planned — for example due to unexpected linac down time — one may want to get as much data as possible on critical targets in the remaining time, regardless of the quality of the data being non-optimal. This sort of tradeoff is very important, and, as will be seen in more detail later, defines the top layer of our model.
@@ 6.1 The Central Decision: What to Measure When and for How Long
However, recall that there may be limited sample material, especially of the most important samples, and the problems may resolve, improving the SNR. These considerations vie in the opposite direction, suggesting that if things are not working well (i.e., for whatever reason, the team is seeing low SNR in secondary data analysis), but it is still early in the overall experiment period, it may be better to hold on to the precious critical sample and hope that the problems resolve. But how long do you wait? Half way? Two thirds? Three quarters? And if the data is marginal, and getting worse, should you continue a measurement to wring as much out of the current sample, or simply abort it and move on to more important samples, given an impending hard stop? And once the ongoing measurement is completed, which of the 8x3x15 minute measurements should be tried next? Perhaps even a previous less-than-satisfactory measurement should be repeated?
@@ 7 A Multi-Scale Cognitive Interaction Model of LCLS Instrument Operations
Our goal in the present work is to predict the potential impacts of UI and workflow changes in the most impactful aspects of LCLS operations. Because of the difficulty of carrying out direct user studies, or of employing a “flight simulator” approach, in the LCLS setting, we chose to create a computer simulation (hereafter, “the model”) of some of the most important human-in-the-loop aspects of experiment operations, and in particular of the phenomena relevant to the core decision-making, as discussed above, of when and whether to continue or abort a measurement, and what to measure next. The goal is to use this model to predict the direction and rough shape of changes in experiment efficiency that might result from changes in instrument procedures and/or control interfaces.
In what follows we discuss the central concepts represented in the model. Much more detail is included in the three video supplements to this paper[vids]. The first video is a concept review, and provides more or less the same information as below. The second includes several demonstrations of the model running, and discusses the results depicted in figures 7 through 9. The third video is a deep dive into the model code, which is available on github.[githubmodel].
@@ 7.2 The Micro-Cognitive Scale: Peak Chasing, Functional Acuity, and Functional Operability
Real experimental decision-making takes place in a very complex setting, where many extrinsic factors are involved, such as which samples are available, what the value is of those samples, the stability and performance of all aspects of the system, how far through the experiment they are, and how much time the remaining high-value samples are likely to take, as well as large-scale factors such as SLAC and LCLS shift-changes and the requirements for relevant members of the team to take breaks to eat and sleep. Our model, of course, represents a highly simplified version of these, but it does not attempt to take into account all aspects of these factors.
@@ 7.3 The Meso-Cognitive Scale: When to End a Measurement
7.4 The Macro-Cognitive Scale: Sample Choice Planning and Re-Planning
The dynamic environment of an LCLS experiment requires a blend of meticulous pre-planning and continuous real-time, opportunistic decision-making and re-planning. This is markedly different from static, puzzle-like problem-solving where one has the luxury of time and a limited set of variables. The actual measurement process is akin to balancing a pencil on its tip, and is made more complicated by the nature of the novel physics being studied at LCLS, and, of course, by a constant barrage of tiny problems, requiring real-time decision-making and re-planning. While some adjustments, like shortening a measurement if one is obtaining high-quality data, are relatively straightforward, others, such as extending a measurement, or even bringing in additional materials from off-site if necessary, can be more challenging given the strictly limited time frame of LCLS experiments, and are potentially impossible. In our model, the only information used to make these decisions is the information available regarding how well the measurements have been going so far, how much beamtime remains, and the number and quality of samples remaining to test.
@@ 7.5 Measurement Error and Performance Quality of Samples
Central to all of the above (and, of course, in the real world LCLS setting) is the quality of the data being gathered. Data quality in settings such as LCLS is often measured in terms of signal-to-noise ratio (SNR), as variously mentioned above. SNR is the slope of a scatter plot of the data, divided by its standard error, and standard error is simply the standard deviation of the data divided by the square root of the number of data points. Therefore, in our model we simply use standard error as a stand-in for (the inverse of) SNR, and the goal of a given measurement is to reach a particular target standard error, which we call “Target Error” (TE), and which is 0.001 by default, for its given sample.
@@ 8 Example Results
Figure 7 depicts two explorations with the model. The left graph in Figure 7 scans across three values of Functional Acuity (FA), and is made with Adjust-Error=False and Cutoff-time=False, thus allowing the user to remain as long as necessary to reach the default Target Error (TE) of 0.001. Recall that runs start at high Performance Quality (PQ) and go to progressively lower PQs — that is, from lower importance samples to higher, or right to left on these graphs. Unsurprisingly, it requires less data to reach the TE threshold (0.001) as PQ increases, and as FA improves (becomes smaller), so that the operator notices that the beam and stream are not aligned at smaller distances of misalignment. The right graph in Figure 7 holds FA to 0.1 (the best performing value from the left), and compares Noticing Delays (NDs). In this case, Adjust-Error=True and Cutoff-time=True, which is the most realistic case: The team is going to get kicked off the machine at a fixed time, but they are allowed to adjust the Error Target to try to get some data on all samples. At first glance, in this case, Noticing Delay makes no apparent difference; All the lines are more or less the same. However, as PQ increases, the amount of data gathered appears to be smaller at both high and low PQ, but higher at middle values of PQ. The reason that this happens is that the experiment manager observes that they are going to run out of time with the default TE, and adjusts the TE upward so that some data is obtained for even the last/most important/lowest PQ samples, although they do run out of time in the end anyway. Again, more flexible re-planning would improve the data quality for the most important (lowest PQ) samples, as will be explored in a separate paper.
@@ 9 Discussion
Modulating Functional Acuity (FA) and Noticing Delay (ND) highlights how improved FA (i.e., faster response to beam misalignment) leads to increased overall experimental efficiency with fewer samples needed for success. In contrast, improved ND produces less significant impacts on efficiency, except in unusual operating regimes of the system, for example when there is no beam time limit. We also investigated quality and duration tradeoffs by adjusting whether the Target Error (our stand in for inverse of SNR) is dynamic or fixed, and by imposing cutoff times, or allowing runs to continue indefinitely. Unsurprisingly, fixed target error requirements produce higher fidelity given unlimited time. However, allowing the target error to adapt yields better time management, ensuring sufficient yields from high performing samples even if lesser quality measurements become truncated by virtue of limited beam time.
@@
Our model merely scratches the surface of the potential of model-based cognitive engineering efforts as applied to scientific user facilities such as LCLS. The present model requires further elaboration and validation with respect to detailed measurements of individual and team behavior. It should be extended to capture more inter- and intra-team communications. Perhaps most importantly, models of learning at all levels should be introduced in order to capture both short term individual learning, for example the instrument scientists becoming familiar with each new experiment; social learning, e.g., the team learning to work together; and longer term learning, such as staff training.
As science probes deeper into the complexities of nature, large scale, beyond-the-state-of-the-art efforts like those at LCLS, and other scientific user facilities, are rapidly increasing in importance. These facilities are in constant flux at every scale. Efforts such as ours that attempt to model the system at multiple scales may be helpful in increasing the efficiency of the science done at these sites by helping to smooth and optimize the human-in-the-loop workflow that will undoubtedly be a significant component of these operations of scientific user facilities for the foreseeable future.
@@ 10 Acknowledgements
The research described here arises from a project begun in 2019 by Devangi Vivrekar and Paul Fuoss, with the assistance of Jeff Shrager. The research was picked up by Teddy Rendahl, and then by Wan-Lin Hu and Jonathan Segal. The numerous LCLS and SLAC staff and visiting scientists that permitted us to observe them working, and sat for interviews over those years, are too numerous to mention, but we are especially grateful to Andy Aquila, Mark Hunter, Meng Liang who welcomed us into their experiments, and to Dan Flath and especially Jana Thayer for valuable discussions and support. We also thank Nick Briggs for a careful reading that uncovered numerous typos. This project was supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences under Contract No. DE-AC02-76SF00515.